OpenStudy (princessaurora):

** WILL FAN & MEDAL ** What is the approximate solution of this system of equations? : y = x^2 + 3x + 3 y = x^2 + x + 2 A) (-0.5, 0.2) B) (-0.5, 1.7) C) (-0.1, 1.2) D) (0.2, 1.5) *I also ask if there is an easier way to solve this type of problem, is there a formula I can use with the calculator?*

OpenStudy (whpalmer4):

If you have a graphing calculator, graph both equations on the same graph, and look for the points where the two cross. Those are the solutions.

OpenStudy (campbell_st):

just equate them \[x^2 + 3x + 3 = x^2 +x+2\] then solve for x

OpenStudy (pawanyadav):

Subtract one equation from another And try to solve the final equation you got after substraction.

OpenStudy (whpalmer4):

@campbell_st has the easiest solution of all for this problem; I'm suggesting ways that you can do this for an arbitrary set of curves that may not be so easily solved this way.

OpenStudy (campbell_st):

when you find x, just substitute it into either equation to find y

OpenStudy (princessaurora):

Thanks a lot that helps!

OpenStudy (whpalmer4):

In a multiple-choice situation like this, you could also evaluate the answer choices and see which one works. For example, if you plug \(x=-0.5 = -1/2\) into the first equation\[x^2+3x+3 = (-1/2)^2+3(-1/2)+3 = 7/4 = 1.6\] and that is not close to the value given in the answer A, so A cannot be the right answer.

OpenStudy (princessaurora):

Oh okay I see

OpenStudy (whpalmer4):

Remember that this problem asks which might be an APPROXIMATE solution, so you can't throw out an answer just because it isn't exactly correct...